OpenStudy (anonymous):

Please help me and not argue on my question, id appreciate it.. Solve x + 2y = 13 4x + 8y = −9 (3, 5) (−1, 7) No Solutions Infinitely Many Solutions

OpenStudy (uri):

x + 2y = 13 4x + 8y = −9 Mutiply the 1st eq with -4 So, -4x-8y=-52 4x + 8y = −9 -------------- WHY YOU TRICK ME???

terenzreignz (terenzreignz):

Trick?

OpenStudy (uri):

Nvm is there a solution to this problem?

terenzreignz (terenzreignz):

Of course not. You can immediately see that they have the same slope. but are not equivalent.

terenzreignz (terenzreignz):

^the only reason I'm saying that so liberally is because you and Jack (I think) have chased the OP out :P

OpenStudy (uri):

So no solution.

terenzreignz (terenzreignz):

Yup.

terenzreignz (terenzreignz):

<takes it back>

OpenStudy (anonymous):

The question is wrongly framed.Not logical since zero cant be equal to 43

terenzreignz (terenzreignz):

It just means there are no solutions @KIEMA ^_^

OpenStudy (anonymous):

Sure.No solutions.What do you think @terenzreignz ?

terenzreignz (terenzreignz):

I certainly think there are no solutions.

OpenStudy (anonymous):

These two equations represent parallel lines in a 2D field. Solutions are those combination where (x,y) satisfy both equations, or more popularly: where these two lines intersect. The intersection point is then the (x,y). As these lines are parallel, they will never intersect, hence there are no solutions. This is option c) 'no solutions' of the question. For info: if these lines would be on top of each other, so the first one would be x+2y=-2 1/4, then every value of (x,y) would fit both equations and in that particular case, answer d) infinite number of solutions would apply.

OpenStudy (unklerhaukus):

The system x + 2y = 13 4x + 8y = −9 if you divide the second equation by four, 4/4x + 8/4y = −9/4 you get x + 2y = -9/4 comparing this to the first equation ? ? 13 = -9/4 ? ?